An association of city mayor conducted a study to determine the average number of times a family went to buy necesarries in a week.They found out that the mean is 4 times in a week

1
Expert's answer
2022-06-07T12:34:24-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le4"

"H_1:\mu>4"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=24" and the critical value for a right-tailed test is "t_c =2.492159."

The rejection region for this right-tailed test is "R = \{t:t>2.492159\}."

The t-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{5-4}{2/\sqrt{25}}=2.5"


Since it is observed that "t=2.5>2.492159=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed, "df=24" degrees of freedom, "t=2.5" is "p=0.009827," and since "p= 0.009827<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 4, at the "\alpha = 0.01" significance level.